OpenStudy (anonymous):

another one :) find the sum and product of the solutions for this equation: (1+i)x^2 +(3-i)x +(2+4i) = 0

OpenStudy (anonymous):

i know that r1r2= c/a and r1 +r2 = -b/a

OpenStudy (anonymous):

By comparing it with: \[ax^2 + bx + c = 0\] Can you tell what is a, b and c here ??

OpenStudy (anonymous):

okay: a= (1+i) b=(3-i) c= (2+4i)

OpenStudy (anonymous):

You have to do same rationalization here..

OpenStudy (anonymous):

ah like radicals?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

thank you, then I can handle it from here :)

OpenStudy (anonymous):

Remember: you will do some mistake in rationalization I guess..

OpenStudy (anonymous):

Be careful: \[i^2 = -1\]

OpenStudy (anonymous):

And sorry not you will do.. I mean to say that you can do some mistake..

OpenStudy (anonymous):

its okay but here is what i did :)

OpenStudy (anonymous):

Yes you can show what you did..

OpenStudy (anonymous):

\[\frac{ -(3-i) }{ (1+i) } \times \frac{ (1-i) }{ (1-i) } \rightarrow \frac{ -3+3i+i-i ^{2} }{ 1-i ^{2} }\]

OpenStudy (anonymous):

so final is : -1+2i

OpenStudy (anonymous):

Yep well done..

OpenStudy (anonymous):

thank you for your help @waterineyes !

OpenStudy (anonymous):

Do the same for product.. Well Done.. Welcome dear..

OpenStudy (anonymous):

yes i did the same for the product and got 3+i =)